### v1

# Active Healing of Microtubule-Motor Networks

Abstract
Cytoskeletal networks have a self-healing property where networks can repair defects to maintain structural integrity. However, both the mechanisms and dynamics of healing remain largely unknown. Here we report an unexplored healing mechanism in microtubule-motor networks by active crosslinking. We directly generate network cracks using a light-controlled microtubule-motor system, and observe that the cracks can self-heal. Combining theory and experiment, we find that the networks must overcome internal elastic resistance in order to heal cracks, giving rise to a bifurcation of dynamics dependent on the initial opening angle of the crack: the crack heals below a critical angle and opens up at larger angles. Simulation of a continuum model reproduces the bifurcation dynamics, revealing the importance of a boundary layer where free motors and microtubules can actively crosslink and thereby heal the crack. We also formulate a simple elastic-rod model that can qualitatively predict the critical angle, which is found to be tunable by two dimensionless geometric parameters, the ratio of the boundary layer and network width, and the aspect ratio of the network. Our results provide a new framework for understanding healing in cytoskeletal networks and designing self-healable biomaterials.
pacs:
Cytoskeletal networks can dynamically reconfigure themselves and generate force to fulfill crucial functions in life, such as mechanical support, motility and division of cells. Furthermore, in neurons, microtubules are organized into parallel arrays that serve as tracks for cargo transport [1]. After an axon is injured, the rearrangement of microtubule orientations into parallel arrays plays a key role in axon regeneration [2]. Kinesin motors that can actively bind and walk on microtubules may contribute to the healing and alignment of microtubule networks [3]. However, such mechanisms of healing by motors have not been studied. Past research on self-healing cytoskeletal networks mainly focuses on mechanisms through adding or reassembling the subunits that make up the cytoskeleton. For example, individual microtubules are found to be capable of incorporating free tubulins to repair lattice defects [4]. At the network level, filamentous actin hydrogels can restore their storage modulus through dynamic polymerization and depolymerization of globular actin, after a shear strain is removed [5]. In this Letter, we investigate an unexplored self-healing mechanism, where active crosslinking by motor proteins can heal cracks in microtubule networks.
Our reconstituted microtubule-motor system [6, 7] provides a light-controllable platform with minimum components that can generate self-healing networks and elucidate the underlying mechanisms. The experimental system [7] consists of free microtubules, light-activatable motor proteins, ATP and buffer solutions, placed in a flow cell, whose height, around 100 $\mu$m, is much smaller than its horizontal dimensions (Supplemental Material). Depolymerization and polymerization of tubulins can be neglected in our experiments. The microtubules are stabilized to minimize depolymerization [6]. Polymerization doubles the average microtubule length, initially around 1.3 $\mu m$, every 4 hours, which is very slow compared to the healing dynamics at the scale of minutes. The engineered motor proteins can “link” under blue light and subsequently crosslink microtubules. Microtubule networks of arbitrary shapes can be generated through light projections onto the flow cell. The networks are contractile due to crosslinking by motors.
We create V-shaped networks to mimic cracks, and find that there exists a critical initial opening angle above which the network buckles, and below which it heals. By healing, we mean the adhesion of two initially separate network arms. Fig. 1(a) and 1(b) show two networks with the same initial arm lengths and widths but different opening angles. The network with the larger angle in Fig. 1(a) keeps opening up as it contracts. Its two arms bend outwards and form a convex shape. In contrast, the network with the smaller initial angle in Fig. 1(b) closes in and the two arms zip up, forming a concave shape. To show healing in Fig. 1(b), we track a small protrusion initially on the right arm, which is labeled by a red circle. The protrusion eventually merges with the left arm (also shown in the Supplemental Movie), demonstrating partial closure of the crack, that is, at least the portion of crack above the protrusion is fully healed. The critical opening angle also depends on the network geometry. We generate two networks with the same arm lengths and opening angles but different widths, as shown in Fig. 1(c) and 1(d). The arm width in Fig. 1(c) is about half of that in Fig. 1(d). As a result, the thinner network buckles outwards while the thicker one bends inwards, indicating that the critical angle can be tuned by the arm shape.
Figure 1: Experiments and simulations of V-shaped active networks show a bifurcation of healing- and buckling-dominated dynamics dependent on the network geometry. (a) and (b) are two networks with the same initial arm lengths and widths but different opening angles. The networks buckle at the large opening angle (a) while heal at the small angle (b). Red circles in (b) track a small protrusion on the right arm which eventually merges with the left arm, demonstrating the partial closure, or healing, of the crack. (c) and (d) are two networks of the same arm lengths and opening angles but different arm widths. The thinner network (c) buckles outwards while the thicker one (d) bends inwards. The spatiotemporal dimensions of the simulated and experimental images are matched. Simulation details are in the Supplemental Material. $t$ ($\overline{t}$) is (dimensionless) time after the first light pulse. In simulations, the microtubule concentration is non-dimensionalized by the initial microtubule concentration. Scale bar, 100 $\mu$m.
The two distinct phenomena in Fig. 1(a) and 1(b) demonstrate a bifurcation of the active network dynamics dependent on the initial opening angles. We denote the dynamics in Fig. 1(a) and Fig. 1(b) the buckling-dominated and healing-dominated regimes, respectively. The two regimes can be differentiated quantitatively by curvature of the network. Given a centerline profile $y(x)$ (inset in Fig. 2), the local curvature $\kappa$ is defined as $\kappa=y^{\prime\prime}/(1+y^{\prime 2})^{3/2}$. We define the mean curvature $[\kappa]$ along the centerline as $[\kappa]=\int_{x_{0}}^{x_{t}}\frac{y^{\prime\prime}}{1+y^{\prime 2}}\mathrm{d}x/\int_{x_{0}}^{x_{t}}\sqrt{1+y^{\prime 2}}\mathrm{d}x$, where $x_{0}$ and $x_{t}$ are the x-coordinates of the starting and ending points of the centerline, respectively. Note that in the definition of $[\kappa]$, the numerator is an integral of the local curvature $\kappa$ over the arc length of the centerline, so it is devided by the total length of the centerline. Time evolutions of mean curvatures in Fig. 1(a) and 1(b) are plotted in Fig. 2. By our definition, a negative curvature represents a concave shape, which emerges in the healing-dominated regime where the attractive healing force dominates and the two network arms caves in to merge, such as Fig. 1(b). In contrast, a positive curvature, that is, a convex shape, arises in the buckling-dominated regime, where the two arms bend outwards and away from each other, as in Fig. 1(a).
Figure 2: Healing and buckling-dominated dynamics can be differentiated by the network curvature. The curvature is negative (concave) for the former and positive (convex) for the latter. The mean curvature $[\kappa]$ is averaged over the centerline of each arm excluding the tip region (dotted lines in the upper right inset), and rescaled by the initial arm length $l$. Error bars represent the difference between left and right arms in a single experiment.
We use simulations based on a field theory model to uncover mechanisms of self-healing and analyze key physical parameters of the network that determine the critical angle in bifurcation. The simulations in Fig. 1 are based on a three-phase model from our previous work [7]: the crosslinked network is modeled as a viscoelastic gel, which self-contracts driven by its internal active stresses; the solvent flow is modeled as a Stokes flow generated by the contraction of the active gel and balanced by the hydrodynamic resistance in the flow cell; the freely-moving microtubules are passive particles (Supplemental Material). The simulation can reproduce the bifurcation dynamics (Fig. 1) and the simulated curvatures are in qualitative agreement with experiments (Fig. 2]). A key step in our simulations is to explicitly model the process of active crosslinking of free microtubules and dimer motors into the crosslinked network, that is
$\frac{\partial\overline{c}}{\partial\overline{t}}+\overline{\bm{\nabla}}\cdot\left(\overline{c}\overline{\bm{v}}\right)=\overline{p}^{\text{on}}_{1}\overline{c}_{f}\overline{d}_{f}-\overline{p}^{\text{off}}_{1}\overline{c},$ (1)
where $\overline{c}$ ($\overline{c}_{f}$) is the crosslinked (free) microtubule concentration, $\overline{d}_{f}$ is the free dimer motor concentration, $\overline{t}$ is time, $\overline{\bm{v}}$ is the crosslinked microtubule velocity, $\overline{p}^{\text{on}}_{1}$ and $\overline{p}^{\text{off}}_{1}$ are the crosslinking and uncrosslinking rates, respectively. Our simulations are non-dimensionalized (Supplemental Material) and the symbols with an overbar represent dimensionless variables. To demonstrate that the self-healing is induced by active crosslinking, we simulate the bifurcation diagram of the network in Fig. 1(d) by varying the crosslinking rates $\overline{p}^{\text{on}}_{1}$ and documenting whether the network buckles or heals under different initial crack angles $\theta$. The result is in Fig. 3, showing that increasing the crosslinking rate renders the network easier to heal and therefore confirming that the healing process is induced by active crosslinking.
Figure 3: Simulated bifurcation diagram shows that the critical angle increases with the crosslinking rate. The diagram documents whether the network in Fig. 1(d) heals or buckles with varying crosslinking rates and initial crack angles. ”Intermediate” represents when the network does not buckle and also does not show significant healing, such as Fig. 1(d), wherein $\overline{p}^{\text{on}}_{1}=15$.
We hypothesize that the microscopic mechanism of active healing is through crosslinking of the free microtubules and motor proteins residing at the boundary of the crosslinked network. The active network is dynamically crosslinked. There are microtubules and motors constantly binding on and falling off the network. We assume there is a layer of free microtubules and dimer motors at the network surface, as shown in Fig. 4. This boundary layer of free microtubules and motors exists due to diffusion and convection of the uncrosslinked components. Consequently, active healing takes place at the overlapped region of the two boundary layers and can bridge the two network arms together (Fig. 4). The boundary-layer model also explains the opening-angle dependency of the bifurcation. The overlapped region expands with decreasing opening angle $\theta$, and therefore healing should prevail at small $\theta$. When the network is healing-dominated, the overlapped region becomes a ”healing front” that can propagate along and zip up the two network arms. The boundary layer of free microtubules and motors can also be visualized in simulations (Fig. S3 in the Supplemental Material).
Figure 4: Schematic of the elastic-rod model. The crosslinked network is modeled as a kinked elastic rod (green). There are boundary layers (light blue) of free motors and microtubules next to the network surfaces. Active crosslinking takes place at the overlapped region (dark blue) of the two boundary layers.
It is well known that an elastic rod under compression can buckle when the compressive load exceeds a threshold. Buckling has also been reported in rectangular active networks [6]. The V-shaped network is a joint of two rectangular networks, where buckling can take place when compression is strong and each arm is long and thin enough. Therefore, we propose that the opening-up dynamics observed in Fig. 1(a) and 1(c) is a result of the buckling instability induced by the compressive active stress. And when buckling takes place, the two arms will only bend outwards in order to reduce the bending energy, which is quadratic in the local curvature [8], at the kink.
Based on the boundary layer and buckling hypotheses, we propose a simple elastic-rod model to predict the critical angle in bifurcation. We assume the crosslinked network is an elastic rod, as shown in Fig. 4. The active stress is assumed to be $\sigma_{a}$, generating a compressive force in each arm that is $F_{a}=\sigma_{a}wd$, with $w$ and $d$ the arm width and depth, respectively. Inside the overlapped boundary layer, active crosslinking introduces an attractive force $F_{h}$ between the two network arms. We assume the surface force density is $\sigma$ (inset in Fig. 4), and the attractive force on each arm is then $F_{h}=\sigma dh/\sin\theta$, where $\theta$ is the opening angle and $h$ is the boundary layer width. The tangential component of $F_{h}$ along the network arm introduces a tension and the minimum compression within each arm is $F_{c}\approx\sigma_{a}wd-\frac{\sigma dh}{2\sin(\theta/2)}$. The bifurcation of network dynamics is governed by two competing mechanisms, one towards buckling, characterized by $F_{c}$, and the other towards healing, characterized by $F_{h}$. We use the classical Euler’s buckling load $F_{c}=\pi^{2}EI/l^{2}$ [9] to approximate the critical load to buckle each network arm, where $E$ is the Young’s modulus, $I=dw^{3}/12$ is the moment of inertia. Therefore, the critical angle $\theta^{*}$ is determined by $F_{c}=F_{b}$, and it follows that
$\theta^{*}=2\arcsin\frac{\sigma h}{2w\left(\sigma_{a}-CE\delta^{2}\right)},$ (2)
where $\delta=w/l$ is the aspect ratio of each arm and $C=\pi^{2}/12$ is a constant. The network will heal when $\theta<\theta^{*}$ and buckle when $\theta>\theta^{*}$. From (2) we can see that $\theta^{*}$ can be tuned by two dimensionless geometric parameters, the ratio of the boundary layer width and the arm width $h/w$, and the aspect ratio of the network $\delta=w/l$. The network shape $w$ and $\delta$ are programmable through light signals in experiments and simulations. Fixing $\delta$ and increasing $w$ will decrease the critical angle $\theta^{*}$. This is because the healing force $F_{h}$ is only determined by $\theta$ and does not depend on $w$ and $\delta$, and the compression $F_{c}$ in each arm is linear in $w$. As the arm width grows, the critical angle needs to shrink in order to generate a greater overlap region, and consequently a larger $F_{h}$, to overcome the increasing compression. Fixing $w$ and decreasing the aspect ratio $\delta$ will also decrease $\theta^{*}$. The compression $F_{c}$ also does not vary with $\delta$ in this case. However, decreasing $\delta$ will make the network more elongated and easier to buckle. Therefore, a smaller critical angle is required to generate a larger attractive force $F_{h}$ in order to overcome the elastic instability.
Figure 5: Bifurcation phase diagrams can be qualitatively predicted by the elastic-rod model. We fix $\delta=0.1$ in (a) and $w=70$ $\mu$m in (b). The fitting parameters used to plot the theoretical curves are $\sigma h/\sigma_{a}=20$ $\mu$m and $\sigma h/CE=1$ $\mu$m.
To test how the network geometry can tune the bifurcation dynamics, we run simulations with different opening angles $\theta$ and vary the network width $w$ and aspect ratio $\delta$. Whether the network buckles or heals in each simulation is documented in the bifurcation diagrams in Fig. 5. We first fix $\delta=0.1$ and simulate networks with different $w$. The bifurcation phase diagram and the theoretical prediction (2) are plotted in Fig. 4(a). As the network width increases, the critical angle $\theta^{*}$ decreases. This confirms our theory (2) that increasing the size of the network while preserving the shape will make buckling easier and healing more difficult. This is because the overlapped boundary layer and the healing force $F_{h}$ do not scale with the network size while the active compression $F_{c}$ is proportional to $w$. As the network dimension increases, the active healing effect becomes less significant. Similarly, the bifurcation phase diagram (Fig. 5(b)) with different $\delta$ and fixed $w$ is also consistent with our theory. As the aspect ratio $\delta$ increases, the network arms become shorter and thus more difficult to buckle. Both simulated phase diagrams are in quantitative agreement with experimental data. In general, $h$, $\sigma$, $\sigma_{a}$ and $E$ in our theory (2) are all functions of the local microtubule and motor concentrations. We treat them as constants in plotting the theory in Fig. 5 for simplicity. Even so, the elastic-rod model can qualitatively predict the bifurcation phase diagram and offers a clear explanation for its dependence on $w$ and $\delta$. Furthermore, we can define two dimensionless groups, $A=2w\sigma_{a}/\sigma h$ and $B=2CE\delta^{2}w/\sigma h$, and rewrite equation (2) as $\theta^{*}=2\arcsin\left[(A-B)^{-1}\right]$, where $A$ is the ratio of the active compression and healing forces, and $B$ is the ratio of the critical buckling load and the healing force. Increasing $A$ or decreasing $B$ renders the network easier to buckle and results in a smaller $\theta^{*}$.
Figure 6: Simulations with different activities $\alpha$ show that the activity does not affect the arm curvatures but can change the time scales and densities of the active networks. Simulation in (a) is the same as Fig. 1(c) and used as a benchmark. Increasing (b) or decreasing (c) the activity $\alpha$ can significantly speed up or slow down the active matter dynamics, respectively. (d) At late times, smaller activity yields a loosely crosslinked network. (e) Depletion of ATP corresponding to (d). The ATP concentration is non-dimensionalized by the initial ATP concentration.
How does activity influence the healing process? The active stress $\bm{\sigma}_{a}$ in our simulations is modeled as an isotropic contractile stress $\sigma_{a}=\alpha c_{A}d_{b}c\bm{I}$, where $\alpha$ is the activity, $c_{A}$ is the ATP concentration, and $d_{b}$ is the concentration of motors bound to the active network (see derivation in the Supplemental Material). We numerically study how the activity $\alpha$ affects the bifurcation dynamics. Using the simulation in Fig. 1(c) as a benchmark, we find that decreasing activity can slow down the active matter dynamics (Fig. 6). The smaller activity also yields a more loosely crosslinked network at late times, as shown in Fig. 6(d), in contrast with the dense networks in Fig. 1(c). However, the activity does not affect the arm curvatures, as seen from Fig. 6(a-c). We also vary the activity on networks with different opening angles and get the same conclusions. The unvarying curvatures indicate that changing activity does not impact the competition of the healing and buckling forces and therefore does not change the critical angle $\theta^{*}$ in bifurcation. This is not surprising as both the healing and compression forces originate from the same microscopic active stress $\bm{\sigma}_{a}$. Our simulations suggest that increasing activity can enhance both healing and compression forces in a similar manner such that even though the overall dynamics is accelerated, the critical angle $\theta^{*}$, which depends on the ratio of the compression and healing forces, remains largely unaffected. The loosely crosslinked network in Fig. 6(d) is a natural consequence of weaker contractile stresses. Additionally, the energy consumption, or equivalently, the ATP consumption also plays an important role in determining the late-stage network densities. In our simulations, the ATP consumption rate is modeled as $-k_{A}d_{b}c_{A}$ with a constant coefficient $k_{A}$ (see equation (S10) in the Supplemental Material). Note that we assume $k_{A}$ is unrelated to the activity $\alpha$. Since a smaller activity prolongs the dynamics and consumes more ATP over a longer time span, at late times the depletion of ATP, as shown in Fig. 6(e), will stall further contraction of the network, which also gives rise to a more loosely crosslinked network.
Cytoskeletal networks provide mechanical support in a wide range of cellular processes and it is important to maintain their integrity under varying loads. Combining experiment and theory, we show that microtubule-motor networks can self-heal through active crosslinking. For V-shaped cracks, the active crosslinking needs to overcome an elastic instability which will open up the crack, leading to a bifurcation of healing and buckling that can be tuned by the initial network geometry. It has been increasingly evident that cytoskeletal networks are gel-like materials [10] and vulnerable to a plethora of mechanical instabilities from self-generated active forces. For example, a bifurcation of in-plane bending and out-of-plane buckling instabilities has been observed in extensile active sheets [11]. Instabilities are not always obstructive. Cells can control and harness instabilities to create useful structures, such as mitotic spindles which are shaped by a barreling-type instability [12]. More work is needed to complete a mechanical-instability phase diagram of active networks and to uncover the regulatory mechanisms used by cells to control such instabilities.
Acknowledgements.
We are grateful to Zhen-Gang Wang and Howard A. Stone for fruitful discussions and to Inna Strazhnik for illustration of Fig. 4. This project is funded by NIH 1R35 GM118043 (MIRA), Packard Foundation, Moore Foundation, and Donna and Benjamin M. Rosen Bioengineering Center. FY also acknowledges financial support from BBE Divisional Postdoctoral Fellowship at Caltech.
Supplemental Material
I Continuum model
In this section we derive a continuum model for the coupled dynamics of active matter and solvent flow. The model details are first published in our previous work [7], where a brief introduction and history of field-theory models on active networks are also included. We reformulate the model below in a more pedagogical manner for readers’ convenience.
I.1 Overview
The main chemical components in our system are microtubules, motors and ATPs. The dynamics of active matter and solvent flows is initiated by a cascade of chemical reactions. Firstly, the monomer motors are activated by light and reversibly dimerized. The dimer motors then crosslink microtubules into a contracting network, which further drives a solvent flow. In our modeling, microtubules and motors are classified according to their chemical structures, such as monomer or dimer motors, and also to their crosslinking status, such as bound dimer motors and free dimer motors, as well as crosslinked and free microtubules. Here ”bound” means the motors are attached to the crosslinked microtubules, and ”free” means the microtubules or motors are not attached to the crosslinked network.
All the chemical reactions mentioned above, namely the reversible dimerization of motors and the crosslinking of microtubules can be modeled in continuity equations of these chemical components. The light patterns enter our model through the dimerization rate $p^{d}(x,y,t)$, where $x$ and $y$ are the spatial coordinates and $t$ is time. The shape of the light region can thereby be determined by a spatiotemporal function $p^{d}(x,y,t)$.
Figure S1: Overview of the three-phase continuum model.
Momentum equations determine the motion of active matter and solvent from their stresses and material properties. In principle, we can write down momentum equations for each chemical species. However, small proteins such as free (monomer and dimer) motors have little mechanical impact on the dynamics and can be assumed to follow the solvent flow. The crosslinked microtubules and bound dimer motors can be modeled together as a single “phase”, which is termed the “crosslinked network”. The crosslinked network is modeled as a viscoelastic gel, which self-contracts driven by its internal active stresses; the solvent flow is modeled as a Stokes flow driven by the contraction of the active gel and balanced by the hydrodynamic resistance in the flow cell; the free microtubules are modeled as passive particles. Each of the three phases exert friction on the other two phases. A summary of the three phases is shown in Fig. S1.
I.2 Continuity Equations
The continuity equations track the change of mass of a component in time due to fluxes and chemical reactions. A general form of continuity equations can be written as
[…]
where the asterisk (*) is used to represent the general form of any chemical species in our system, $t$ is time, $c^{*}$ is concentration, and $\bm{j}^{*}$ is flux. In our system, the flux $\bm{j}^{*}$ usually consist of two components, i.e.,
$\bm{j}^{*}=c^{*}\bm{v}^{*}-D^{*}\bm{\nabla}c^{*},$ (S2)
where $\bm{v}^{*}$ is velocity and $D^{*}$ is diffusivity. $c^{*}\bm{v}^{*}$ and $-D^{*}\bm{\nabla}c^{*}$ are called the convective and diffusive fluxes, respectively. We follow a convention to keep the convective flux on the left-hand side of (S1) while moving the diffusive flux to the right-hand side. Therefore, the general form of continuity equations is
$\frac{\partial c^{*}}{\partial t}+\bm{\nabla}\cdot\left(c^{*}\bm{v}^{*}\right)=\{\mbox{chemical-reaction terms}\}+D^{*}\nabla^{2}c^{*}$ (S3)
Figure S2: Illustrations of chemical reactions in the model. A, Reversible dimerization of motors. B, Reversible crosslinking of a freely-moving microtubule $i$ and another microtubule $j$. The latter can be either free or pre-crosslinked. C, Reversible binding of a dimer motor on two pre-crosslinked microtubules.
The chemical reactions in our system are sketched in Fig. S2. The motor proteins can reversibly dimerize and the continuity equation is
$\frac{\partial m}{\partial t}+\bm{\nabla}\cdot(m\bm{u})=2(p^{m}d_{f}-p^{d}m^{2})+D_{m}\nabla^{2}m$ (S4)
where $m$ is the monomer motor concentration, $d_{f}$ is the free dimer concentration, $\bm{u}$ is the solvent flow velocity, $D_{m}$ is the diffusivity, $p^{m}$ and $p^{d}(x,y,t)$ are the monomerization and dimerization rates, respectively. The function $p^{d}(x,y,t)$ is determined by the light pattern. In experiments, we use V-shaped light patterns, which can be defined by a set of points $A_{l}=\{(x,y)\in\mbox{light pattern}\}$. The light pattern is projected periodically onto the flow cell with time interval $T$ and duration $\Delta t$. Naively, the dimerization rate can be written as
$p^{d}_{i}(x,y,t)=\begin{cases}p^{d}_{0}&(x,y)\in A_{l}\ \mbox{and }t\in[nT,nT+\Delta t],\\ 0&\text{otherwise},\end{cases}$ (S5)
where $p^{d}_{0}$ is a constant dimerization rate and $n=0,1,2,\dots$ is the number of light pulses. However, directly using (S5) in simulations will cause numerical instability due to discontinuity of $p^{d}_{i}(x,y,t)$ in space. There are multiple ways to deal with this. In this Letter, we use a Gaussian kernel smoother to make (S5) a smooth function in space. The Gaussian kernel is defined as
$K_{r}(x,y,X,Y)=e^{-\frac{(x-X)^{2}+(y-Y)^{2}}{2r^{2}}}$ (S6)
with a constant $r$, which is the “kernel radius”. Denoting the whole calculation domain as $A$, the smooth dimerization rate $p^{d}(x,y,t)$ can be obtained from applying the smoother (S6) on (S5) through
$p^{d}(x,y,t)=\frac{\int_{A}p^{d}_{i}(X,Y,t)K_{r}(x,y,X,Y)\mathrm{d}X\mathrm{d}Y}{\int_{A}K_{r}(x,y,X,Y)\mathrm{d}X\mathrm{d}Y}.$ (S7)
The slope of $p^{d}(x,y,t)$ at the boundary of light pattern $A_{l}$ can be tuned by the parameter $r$. A smaller $r$ yields a larger slope and thereby a sharper light boundary in simulations.
The continuity equations for the crosslinked ($c$) and free ($c_{f}$) microtubules are
$\frac{\partial c}{\partial t}+\bm{\nabla}\cdot\left(c\bm{v}\right)=p^{\text{on}}_{1}c_{f}d_{f}-p^{\text{off}}_{1}c,\\ \frac{\partial c_{f}}{\partial t}+\bm{\nabla}\cdot\left(c_{f}\bm{v}_{f}\right)=-p^{\text{on}}_{1}c_{f}d_{f}+p^{\text{off}}_{1}c+D_{f}\nabla^{2}c_{f}.$ (S8)
where $D_{f}$ is the diffusivity of the free microtubules, $p^{\text{on}}_{1}$ and $p^{\text{on}}_{1}$ are the crosslinking and un-crosslinking rates, respectively, see Fig. S2B. The velocities of the crosslinked and free microtubules are denoted by $\bm{v}$ and $\bm{v}_{f}$, respectively. In (S8) we neglect the diffusion flux of the crosslinked microtubules. Also Note that in writing down the chemical reaction terms in (S8), we assume the reaction orders are 1 for simplicity.
The continuity equations for the free ($d_{f}$) and the bound ($d_{b}$) dimer motors are
$\frac{\partial d_{f}}{\partial t}+\bm{\nabla}\cdot\left(d_{f}\bm{u}\right)=-p^{m}d_{f}+p^{d}m^{2}-p_{1}^{\text{on}}c_{f}d_{f}-p_{2}^{\text{on}}cd_{f}+p^{\text{off}}_{2}d_{b}+D_{d}\nabla^{2}d_{f}\\ \frac{\partial d_{b}}{\partial t}+\bm{\nabla}\cdot\left(d_{b}\bm{v}\right)=p_{1}^{\text{on}}c_{f}d_{f}+p_{2}^{\text{on}}cd_{f}-p^{\text{off}}_{2}d_{b},$ (S9)
where $d_{b}$ is the bound dimer concentration, $p_{2}^{\text{on}}$ is the rate of the dimer motors binding on the crosslinked microtubules, $D_{d}$ is the diffusivity of the free dimers, $p^{\text{off}}_{2}$ is the unbinding rate of a dimer motor that does not uncrosslink the two microtubules, see Fig. S2C. All free motors are assumed to follow the solvent flow ($\bm{u}$) and the bound motors follow the crosslinked gel ($\bm{v}$) .
The continuity equation for the ATP is
$\frac{\partial c_{A}}{\partial t}+\bm{\nabla}\cdot(c_{A}\bm{u})=-k_{A}d_{b}c_{A}+D_{A}\nabla^{2}c_{A},$ (S10)
where $k_{A}$ is the consumption rate of the ATP by the bound dimer motors, and $D_{A}$ is the ATP diffusivity.
The solvent flow is incompressible, meaning that its density does not change. The continuity equation for the solvent flow is just
$\bm{\nabla}\cdot\bm{u}=0.$ (S11)
I.3 Momentum Equations
Momentum equations are the balance between the momentum fluxes, stresses and body forces in a differential form. At low Reynolds number, inertial momentum fluxes are usually neglected. Under this assumption, a general form of the momentum equation is
$\bm{\nabla}\cdot\bm{\sigma}^{*}+\bm{f}^{*}=\bm{0},$ (S12)
where $\bm{\sigma}^{*}$ is the total stress and $\bm{f}^{*}$ is the total body force. In this section, we list the momentum equations for each phase.
I.3.1 Crosslinked network as an active gel
The crosslinked network forms an active gel, which is assumed to be viscoelastic. The momentum equation can be written as
$\bm{\nabla}\cdot\left(\bm{\sigma}_{a}+\bm{\sigma}_{v}+\bm{\sigma}_{el}+\bm{\sigma}_{st}\right)+\bm{f}_{fl}+\bm{f}_{f}=\bm{0},$ (S13)
where $\bm{\sigma}_{a}$, $\bm{\sigma}_{v}$, $\bm{\sigma}_{el}$ and $\bm{\sigma}_{st}$ are the active, viscous, elastic and steric stresses in the gel, $\bm{f}_{fl}$ and $\bm{f}_{f}$ are the friction with the solvent and freely-moving microtubules, respectively. Their explicit expressions can be derived from microscopic interactions [13, 14]. The active contractile stress is assumed to be
$\bm{\sigma}_{a}=\alpha c_{A}d_{b}c\bm{I},$ (S14)
where $c$, $c_{A}$ and $d_{b}$ are the concentrations of the crosslinked microtubules, ATP and the bound dimer motors, respectively, $\alpha$ is a constant activity coefficient and $\bm{I}$ is the identity tensor. This active stress is derived to be $\bm{\sigma}^{\prime}_{a}\propto c^{2}\bm{I},$ in Ref [14], where the bound motor concentration is implicitly assumed to be uniform and does not appear in active stresses. In our experiments, the concentration difference of the dimer motors inside and outside of the illuminated regions is the key to form active gels. We can modify the theory in Ref [14] by adding a prefactor $d_{b}/c$ in the active stresses, which is the number of bound motors per microtubule.
The viscous stress is
$\bm{\sigma}_{v}=\eta d_{b}c\left(\bm{\nabla}\bm{v}+\bm{\nabla}\bm{v}^{T}\right),$ (S15)
where $\bm{v}$ is the velocity of the crosslinked microtubules and $\eta$ is the viscosity coefficient. This viscous stress in Ref [14] is $\bm{\sigma}^{\prime}_{v}\propto c^{2}\left(\bm{\nabla}\bm{v}+\bm{\nabla}\bm{v}^{T}\right)$. In our model the coefficient is modified from $c^{2}$ to $d_{b}c$ to incorporate the effects of the motor concentrations.
The elastic stress $\bm{\sigma}_{el}$ is assumed to follow the Oldroyd-Maxwell model with a long relaxation time [15],
$\frac{\mathrm{D}\bm{\sigma}_{el}}{\mathrm{D}t}-\left[\bm{\sigma}_{el}\cdot\bm{\nabla}\bm{v}+\bm{\nabla}\bm{v}^{T}\cdot\bm{\sigma}_{el}\right]=\eta_{el}d_{b}c\left(\bm{\nabla}\bm{v}+\bm{\nabla}\bm{v}^{T}\right)-p^{\text{off}}_{1}\bm{\sigma}_{el},$ (S16)
where $\eta_{el}$ is the elasticity coefficient and $\mathrm{D}/\mathrm{D}t$ is the material derivative. We choose the Oldroyd-Maxwell model for its simplicity.
To avoid unlimited density build-up at one point, we add a steric stress that is
$\bm{\sigma}_{st}=-\xi c^{2}\bm{I},$ (S17)
with a constant coefficient $\xi$.
The friction with freely-moving microtubules is
$\bm{f}_{f}=\beta cc_{f}(\bm{v}_{f}-\bm{v}),$ (S18)
where $\beta$ is the friction coefficient. The friction with the solvent flow is
$\bm{f}_{fl}=\gamma c(\bm{u}-\bm{v})$ (S19)
with the drag coefficient $\gamma$. In general the hydrodynamic drag coefficient depends on the fiber orientation [16]. In this Letter we neglect the polarity in both active stresses and hydrodynamic friction.
I.3.2 Free microtubules
The free microtubules are passive particles that only experience friction from the gel and the ambient fluid. The force balance is $\beta c_{f}c\left(\bm{v}-\bm{v}_{f}\right)+\gamma c_{f}(\bm{u}-\bm{v}_{f})=\bm{0}$, and their velocity is therefore
$\bm{v}_{f}=\frac{\beta c\bm{v}+\gamma\bm{u}}{\beta c+\gamma}.$ (S20)
I.3.3 Solvent flows
The flow cells used in our experiments are usually called Hele-Shaw cells in fluid mechanics, meaning that their horizontal dimensions ($x-$ and $y-$direction) greatly exceed their vertical dimension ($z$-direction). We now derive the averaged two-dimensional flow equation in the $xy-$plane. We start from the three-dimensional Stokes equation, that is,
~$-\widetilde{\bm{\nabla}}\widetilde{\Pi}+\mu\widetilde{\nabla}^{2}\widetilde{\bm{{u}}}+\gamma\widetilde{c}\left(\widetilde{\bm{{v}}}-\widetilde{\bm{{u}}}\right)+\gamma\widetilde{c}_{f}\left(\widetilde{\bm{{v}}}_{f}-\widetilde{\bm{{u}}}\right)=0,$ (S21)
where $\widetilde{\Pi}$ and $\mu$ are the fluid pressure and viscosity, respectively. We use a tilde $\sim$” on top to indicate the variable is a function of $(x,y,z)$ and if the variable is a vector, it is a three-dimensional vector, e.g., $\widetilde{\Pi}=\widetilde{\Pi}(x,y,z)$ and $\widetilde{\bm{{u}}}(x,y,z)=\left(\widetilde{u}_{x}(x,y,z),\widetilde{u}_{y}(x,y,z),\widetilde{u}_{z}(x,y,z)\right)$. Since the vertical length scale is much smaller than horizontal length scales, the classical lubrication theory [16] shows that $\widetilde{\Pi}(x,y,z)\approx\Pi(x,y)$, i.e., the fluid pressure is constant along the $z$-direction, and $\widetilde{u}_{z}\approx 0$. Substituting these two results into equation (S21) yields $\widetilde{v}_{z}\approx 0$ and $\widetilde{v}_{f,z}\approx 0$. Furthermore, the $x-$ and $y-$components of equation (S21) can be approximated by
$-\bm{\nabla}\Pi+\mu\frac{\partial^{2}\widetilde{\bm{{u}}}}{\partial z^{2}}+\gamma\widetilde{c}\left(\widetilde{\bm{{v}}}-\widetilde{\bm{{u}}}\right)+\gamma\widetilde{c}_{f}\left(\widetilde{\bm{{v}}}_{f}-\widetilde{\bm{{u}}}\right)=0,$ (S22)
where all vectors only have $x-$ and $y-$components, i.e., $\widetilde{\bm{{u}}}=\left(\widetilde{u}_{x}(x,y,z),\widetilde{u}_{y}(x,y,z)\right)$, $\widetilde{\bm{{v}}}=\left(\widetilde{v}_{x}(x,y,z),\widetilde{v}_{y}(x,y,z)\right)$ and $\widetilde{\bm{{v}}}_{f}=\left(\widetilde{v}_{f,x}(x,y,z),\widetilde{v}_{f,y}(x,y,z)\right)$. Therefore, the analytical solution of $\widetilde{\bm{{u}}}$ requires knowledge of the three-dimensional distribution of $\widetilde{c}$, $\widetilde{\bm{{v}}}$ and $\widetilde{\bm{{v}}}_{f}$. To derive a two-dimensional model, we further simplify equation (S22) by replacing the friction with their $z$-directional average, i.e.,
$-\bm{\nabla}\Pi+\mu\frac{\partial^{2}\widetilde{\bm{{u}}}}{\partial z^{2}}+\gamma c\left(\bm{v}-\bm{u}\right)+\gamma c_{f}\left(\bm{v}_{f}-\bm{u}\right)=0,$ (S23)
where $c(x,y)=h^{-1}\int_{0}^{h}\widetilde{c}\mathrm{d}z$, $\bm{v}(x,y)=h^{-1}\int_{0}^{h}\widetilde{\bm{{v}}}\mathrm{d}z$, $\bm{v}_{f}(x,y)=h^{-1}\int_{0}^{h}\widetilde{\bm{{v}}}_{f}\mathrm{d}z$, and $\bm{u}(x,y)=h^{-1}\int_{0}^{h}\widetilde{\bm{{u}}}\mathrm{d}z$ with the top and bottom walls of channel at $z=0$ and $z=h$, respectively. These notations are consistent with the rest of the Letter. The solution of equation (S23) is
$\widetilde{\bm{{u}}}=\frac{1}{2\mu}\left[\bm{\nabla}\Pi-\gamma c\left(\bm{v}-\bm{u}\right)-\gamma c_{f}\left(\bm{v}_{f}-\bm{u}\right)\right]z(z-h).$ (S24)
Combining $\bm{u}=h^{-1}\int_{0}^{h}\widetilde{\bm{{u}}}\mathrm{d}z$ and equations (S20) and (S24), we have the equation for the averaged two-dimensional flow velocity
$\bm{u}=-\frac{h^{2}}{12\mu}\left[\bm{\nabla}\Pi-\left(\gamma c+\gamma_{f}c_{f}\right)\left(\bm{v}-\bm{u}\right)\right],$ (S25)
with $\gamma_{f}=\gamma\beta c/(\beta c+\gamma)$. The above result is a modified Darcy’s law in a Hele-Shaw cell [16] by incorporating the microtubule friction.
I.4 Non-dimensionalization
Using the initial microtubule concentration $c_{0}$, initial monomer motor concentration $m_{0}$, initial ATP concentration $c_{A0}$, and the typical length scale of the illuminated region $l$, we can non-dimensionalize the governing equations by
$\begin{split}\overline{c}=\frac{c}{c_{0}},\quad\overline{\bm{{\nabla}}}=l_{0}\bm{\nabla},\quad\overline{m}=\frac{m}{m_{0}},\quad\overline{d}_{f}=\frac{d_{f}}{m_{0}},\quad\overline{d}_{b}=\frac{d_{b}}{m_{0}},\quad\overline{c}_{A}=\frac{c_{A}}{c_{A0}}\\ \overline{t}=\frac{t}{t_{0}},\quad\overline{\bm{{v}}}=\frac{\bm{v}}{v_{0}},\quad\overline{\bm{{v}}}_{f}=\frac{\bm{v}_{f}}{v_{0}},\quad\overline{\bm{{u}}}=\frac{\bm{u}}{v_{0}},\quad\overline{\bm{{\sigma}}}_{el}=\frac{\bm{\sigma}_{el}}{\sigma_{0}},\quad\overline{\Pi}=\frac{\Pi}{\Pi_{0}},\end{split}$ (S26)
where we use overlines to denote dimensionless variables, $t_{0}$, $v_{0}$, $\sigma_{0}$ and $\Pi_{0}$ are typical scales of time, velocity, stress, and fluid pressure, respectively. By balancing the contractile stress (S14) and the viscous stress (S15) in the active gel, we have $v_{0}=\alpha c_{A0}l/\eta$. The typical time scale follows as $t_{0}=l_{0}/v_{0}=\eta/\alpha c_{A0}$. The typical stress scale can be obtained from equation (S14), which is $\sigma_{0}=\alpha c_{A0}c_{0}m_{0}$. From equation (S25), we have $\Pi_{0}=c_{0}\gamma v_{0}l_{0}=\alpha\gamma c_{0}c_{A0}l_{0}^{2}/\eta$.
The dimensionless continuity equations for the monomer (S4) and dimer (S9) motors are
$\frac{\partial\overline{m}}{\partial\overline{t}}+\overline{\bm{{\nabla}}}\cdot(\overline{m}\overline{\bm{{u}}})=2(\overline{p}^{m}\overline{d}_{f}-\overline{p}^{d}\overline{m}^{2})+\text{Pe}_{m}^{-1}\overline{\nabla}^{2}\overline{m},\\ \frac{\partial\overline{d}_{f}}{\partial\overline{t}}+\overline{\bm{{\nabla}}}\cdot\left(\overline{d}_{f}\overline{\bm{{u}}}\right)=-\overline{p}^{m}\overline{d}_{f}+\overline{p}^{d}\overline{m}^{2}-\overline{p}_{1}^{\text{on}}\overline{c}_{f}\overline{d}_{f}-\overline{p}_{2}^{\text{on}}\overline{c}\overline{d}_{f}+\overline{p}^{\text{off}}_{2}\overline{d}_{b}+\text{Pe}_{d}^{-1}\overline{\nabla}^{2}\overline{d}_{f},\\ \frac{\partial\overline{d}_{b}}{\partial\overline{t}}+\overline{\bm{{\nabla}}}\cdot\left(\overline{d}_{b}\overline{\bm{{v}}}\right)=\overline{p}_{1}^{\text{on}}\overline{c}_{f}\overline{d}_{f}+\overline{p}_{2}^{\text{on}}\overline{c}\overline{d}_{f}-\overline{p}^{\text{off}}_{2}\overline{d}_{b},$ (S27)
where the dimensionless reaction coefficients are $\overline{p}^{m}=p^{m}t_{0}$, $\overline{p}^{d}=p^{m}m_{0}t_{0}$, $\overline{p}^{\text{on}}_{1}=p^{\text{on}}_{1}c_{0}t_{0}$, $\overline{p}^{\text{on}}_{2}=p^{\text{on}}_{2}c_{0}t_{0}$ and $\overline{p}^{\text{off}}_{2}=p^{\text{off}}_{2}t_{0}$. We use Pe to denote the Péclet number and $\text{Pe}_{m}=v_{0}l/D_{m}$, $\text{Pe}_{d}=v_{0}l/D_{d}$.
The dimensionless continuity equations for the microtubules (S8) are
$\frac{\partial\overline{c}}{\partial\overline{t}}+\overline{\bm{{\nabla}}}\cdot\left(\overline{c}\overline{\bm{{v}}}\right)=\overline{p}^{\text{on}}_{1}\overline{c}_{f}\overline{d}_{f}-\overline{p}^{\text{off}}_{1}\overline{c},\\ \frac{\partial\overline{c}_{f}}{\partial\overline{t}}+\overline{\bm{{\nabla}}}\cdot\left(\overline{c}_{f}\overline{\bm{{v}}}_{f}\right)=-\overline{p}^{\text{on}}_{1}\overline{c}_{f}\overline{d}_{f}+\overline{p}^{\text{off}}_{1}\overline{c}+\text{Pe}_{f}^{-1}\overline{\nabla}^{2}\overline{c}_{f}$ (S28)
with $\overline{p}^{\text{off}}_{1}=p^{\text{off}}_{1}t_{0}$, $\text{Pe}_{f}=v_{0}l_{0}/D_{f}$. For the ATP, equation (S10) becomes
$\frac{\partial\overline{c}_{A}}{\partial\overline{t}}+\overline{\bm{{\nabla}}}\cdot\left(\overline{c}\overline{\bm{{u}}}\right)=-\overline{k}_{A}\overline{d}_{b}\overline{c}_{A}+Pe_{A}^{-1}\overline{\nabla}^{2}\overline{c}_{A},$ (S29)
where $\overline{k}_{A}=k_{A}m_{0}t_{0}$ and $\text{Pe}_{A}=v_{0}l_{0}/D_{A}$.
The continuity equation for the solvent flow (S11) is
¯$\overline{\bm{{\nabla}}}\cdot\overline{\bm{{u}}}=0.$ (S30)
Using equations (S13-S20), the dimensionless momentum equations for the active gel are
¯$\overline{\bm{{\nabla}}}\cdot\left[\left(\overline{c}_{A}\overline{d}_{b}\overline{c}-\overline{\xi}\overline{c}^{2}\right)\bm{I}+\overline{\bm{{\nabla}}}\overline{\bm{{v}}}+\overline{\bm{{\nabla}}}\overline{\bm{{v}}}^{T}+\overline{\bm{{\sigma}}}_{el}\right]+\left(\overline{\gamma}_{f}\overline{c}_{f}+\overline{\gamma}\overline{c}\right)\left(\overline{\bm{{u}}}-\overline{\bm{{v}}}\right)=\bm{0}$ (S31)
where $\overline{\xi}=\xi c_{0}/\alpha c_{A0}m_{0}$, $\overline{\gamma}=\gamma l_{0}^{2}/m_{0}\eta$, $\overline{\gamma}_{f}=\gamma_{f}l_{0}^{2}/m_{0}\eta$ and
$\frac{\mathrm{D}\overline{\bm{{\sigma}}}_{el}}{\mathrm{D}\overline{t}}-\left[\overline{\bm{{\sigma}}}_{el}\cdot\overline{\bm{{\nabla}}}\overline{\bm{{v}}}+\overline{\bm{{\nabla}}}\overline{\bm{{v}}}^{T}\cdot\overline{\bm{{\sigma}}}_{el}\right]=\overline{\eta}_{el}\overline{d}_{b}\overline{c}\left(\overline{\bm{{\nabla}}}\overline{\bm{{v}}}+\overline{\bm{{\nabla}}}\overline{\bm{{v}}}^{T}\right)-\overline{p}^{\text{off}}_{1}\overline{\bm{{\sigma}}}_{el},$ (S32)
with $\overline{\eta}_{el}=\eta_{el}/\alpha c_{A0}$.
The dimensionless velocity of the freely-moving microtubules (S20) is
$\overline{\bm{{v}}}_{f}=\frac{\overline{\beta}\overline{c}\overline{\bm{{v}}}+\overline{\gamma}\overline{\bm{{u}}}}{\overline{\beta}\overline{c}+\overline{\gamma}},$ (S33)
with $\overline{\beta}=\beta l_{0}^{2}c_{0}/m_{0}\eta$. Note that only two of $\overline{\beta}$, $\overline{\gamma}$ and $\overline{\gamma}_{f}$ are independent, which are connected through $\overline{\gamma}_{f}=\overline{\gamma}\overline{\beta}\overline{c}/(\overline{\beta}\overline{c}+\overline{\gamma})$.
The dimensionless solvent flow (S25) is
$\overline{\bm{{u}}}=-\overline{\zeta}\left[\overline{\bm{{\nabla}}}\overline{\Pi}-\left(\overline{c}+\frac{\overline{\gamma}_{f}}{\overline{\gamma}}\overline{c}_{f}\right)\left(\overline{\bm{{v}}}-\overline{\bm{{u}}}\right)\right],$ (S34)
with $\overline{\zeta}=h^{2}c_{0}\gamma/12\mu$.
I.5 Simulations
Figure S3: Normalized concentration contours of crosslinked microtubules, free microtubules and free dimer motors. The contours of crosslinked microtubules (a) are encapsulated by the contours of of free microtubules (b) and free dimer motors (c) at the same levels, demonstrating that there is a boundary layer of free microtubules and motors outside the network surface. (e) and (f) are overlaid contours of crosslinked and free microtubules, and crosslinked and free dimer motors, respectively.
We used the finite difference method in numerical simulations with the central difference scheme in space and the method of lines in time. The codes are written in Python and available at https://github.com/fy26/ActiveHealing.
The parameters used in simulations are
$l_{0}=220\ \mu\mbox{m},\quad t_{0}=40\ \mbox{s},\quad p^{d}_{0}=15,\quad p^{m}=12,\quad p^{\text{on}}_{1}=15,\quad p^{\text{on}}_{2}=0.75,$
$p^{\text{off}}_{1}=10,\quad p^{\text{off}}_{2}=20,\quad\text{Pe}_{f}^{-1}=10^{-3},\quad\text{Pe}_{m}^{-1}=\text{Pe}_{d}^{-1}=\text{Pe}_{A}^{-1}=10^{-2},$
$\overline{k}_{A}=0.8,\quad\overline{\xi}=0.08,\quad\overline{\eta}_{el}=1.5,\quad\overline{\gamma}=0.045,\quad\overline{\beta}=0.45,\quad\overline{\zeta}=20.5.$ (S35)
In addition, the duration of each light pulse, $\Delta t/t_{0}=0.04$ and the time interval between two light pulses is $T/t_{0}=0.2$.
I.5.1 Boundary layer of free microtubules and motors
We plot normalized concentrations of crosslinked microtubules, free microtubules and free dimer motors in Fig. S3. The contours are simulated results of Fig 1(b) in the main text at $\overline{t}=3$. From Fig. S3 we can see that the contours of crosslinked microtubules are always encapsulated by the contours of free microtubules and motors, confirming that there is a boundary layer at the surface of active networks.
II Experimental Setup
II.1 Light-induced kinesin expression and purification
We constructed two chimeras of D. melanogaster kinesin K401: K401-iLID and K401-micro as previously described (Addgene 122484 and 122485) [6]. In short, for the K401-iLiD plasmid we inserted iLID with a His tag after the C-terminus of K401. For the K401-micro plasmid, we inserted K401 between the His-MBP and the micro. The MBP domain is needed to ensure the microdomain remains fully functional during expression [17]. After the expression, the MBP domain can be cleaved of by TEV protease.
For protein expression, we transformed the plasmids to BL21(DE3)pLysS cells. The cells were grown in LB and induced at with 1mM IPTG at 18C for 16 hours after reaching OD 0.6. The cells were then pelleted at 4000G and resuspended in lysis buffer (50 mM sodium phosphate, 4 mM MgCl2, 250 mM NaCl, 25 mM imidazole, 0.05 mM MgATP, 5 mM BME, 1 mg/ml lysozyme and 1 tablet/50 mL of Complete Protease Inhibitor). After a 1-hour incubation with stirring, the lysate was passed through a 30kPSI cell disruptor. The lysate was clarified at 30,000 G for 1 hour. The supernatant was then incubated with Ni-NTA agarose resin for 1 hour. The lysate/Ni-NTA mixture was loaded into a chromatography column and washed three times with wash buffer (Lysis buffer with no lysozyme nor complete EDTA tablet), and eluted with 500mM imidazole. Protein elutions were dialyzed overnight using 30 kDa MWCO membrane against 50 mM sodium phosphase, 4 mM MgCl2, 250 mM NaCl, 0.05 mM MgATP, and 1 mM BME. For the K401-micro elution, we added TEV protease at a 1:25 mass ratio to remove the MBP domain. Then, we used centrifugal filters to exchange to pH 6.7 protein storage buffer (50 mM imidazole, HCl for pH balancing, 4 mM MgCl2, 2 mM DTT, 50 $\mu$M MgATP, and 36% sucrose). Proteins were then aliquoted and flash frozen in LN2 and stored under -80C.
II.2 Microtubule Polymerization and Length Distribution
Fluorscent microtubule polymerization was previously described [6]. In short, we used a protocol based on one found on the Dogic lab homepage. The procedure began by preloading and starting a 37C water bath. GMP-cpp, reagents, and tubes were cooled on ice. A 20 mM DTT solution was prepared using Pierce no-weigh format, and a GMP mixture consisting of M2B, DTT, and GMP-cpp was made and stored on ice. The ultracentrifuge and rotor were pre-cooled to 4C. Tubulin (20 mg/mL) and labeled tubulin (20 mg/mL) were thawed in the water bath until mostly thawed and then cooled on ice. In a cold room, labeled tubulin was added to the stock vial of unlabeled tubulin and mixed gently. The GMP mixture was then added to the tubulin mixture and stirred gently. This combined mixture was pipetted into ultracentrifuge tubes and incubated on ice for 5 minutes before being centrifuged at 90,000 rpm, 4C for 8 minutes. The supernatant was carefully collected without disturbing the pellet and transferred to an Eppendorf tube, mixed, and stored on ice. The mixture was incubated in a 37C water bath for 1 hour, protected from light. Aliquots were then dispensed into PCR strip tubes, which were spun to collect the fluid at the bottom. Finally, the PCR strips were flash frozen in liquid nitrogen and stored in a -80C freezer.
To measure the length distribution of microtubules, we imaged fluorescently labeled microtubules immobilized onto the cover glass surface of a flow cell. The cover glass was treated with a 0.01% solution of poly-L-lysine (Sigma P4707) to promote microtubule binding. The lengths of microtubules were determined by image segmentation. Each microtubule image was normalized and underwent local and global thresholding to correct for non-uniform backgrounds and obtain thresholded images of putative microtubules. Morphological operations were applied to reconnect small breaks in filaments. Objects near the image boundary were removed, and small or circular objects were filtered out based on size and eccentricity thresholds. Potential microtubule crossovers were identified and eliminated by analyzing the angles of lines within the image.
II.3 Sample Chambers for Experiments
Glass slides and coverslips were first cleaned using a series of washes. Slides and coverslips were placed in respective containers, and 2% Hellmanex solution was prepared by mixing 6 mL Hellmanex with 300 mL DI water, heated, and poured into the containers. The containers were sonicated for 10 minutes, followed by three DI water rinses and an ethanol rinse. Ethanol was added to the containers and sonicated again for 10 minutes, followed by another ethanol rinse and three DI water rinses. Next, 0.1 M KOH was added to the containers, sonicated for 10 minutes, and rinsed three times with DI water. The slides were then etched overnight with 5% HCl and rinsed three times with DI water. Clean slides and coverslips were stored in DI water.
For silane coupling, a 2% acrylamide solution was prepared using 40% acrylamide stock solution, and degassed under vacuum. In a chemical hood, 98.5% ethanol, 1% acetic acid, and 0.5% silane agent were mixed to prepare the silane-coupling solution, which was immediately poured into the containers with the slides and coverslips and incubated at room temperature for 20-30 minutes. The slides were then rinsed once with ethanol, three times with DI water, and baked at 110C for 30 minutes or 50C overnight.
For acrylamide polymerization, the degassed 300 mL of 2% acrylamide solution was moved to a stir plate, and 105 $\mu$L TEMED and 210 mg ammonium persulfate were added. The solution was immediately poured over the silane-coupled slides and coverslips and left to polymerize overnight at 4C. Before use, the slides and coverslips were rinsed with DI water and air-dried.
II.4 Reaction Mixture
For the self-organization experiments, K401-micro, K401-iLID, and microtubules were combined into a reaction mixture to achieve final concentrations of approximately 0.1 $\mu$M for each motor type and 1.5-2.5 $\mu$M for tubulin, referring to protein monomers for K401-micro and K401-iLID constructs, and protein dimers for tubulin. The sample preparation was conducted under dark-room conditions to minimize unintended light activation, using room light filtered to block wavelengths below 580 nm (Kodak Wratten Filter No. 25). The base reaction mixture included a buffer, MgATP as an energy source, glycerol as a crowding agent, pluronic F-127 for surface passivation, and components for oxygen scavenging (pyranose oxidase, glucose, catalase, Trolox, DTT), along with ATP-recycling reagents (pyruvate kinase/lactic dehydrogenase, phosphoenolpyruvic acid). The reaction mixture consisted of 59.2 mM K-PIPES pH 6.8, 4.7 mM MgCl2, 3.2 mM potassium chloride, 2.6 mM potassium phosphate, 0.74 mM EGTA, 1.4 mM MgATP, 10% glycerol, 0.50 mg/mL pluronic F-127, 2.9mg/mL pyranose oxidase, 3.2 mg/mL glucose, 0.086 mg/mL catalase, 5.4 mM DTT, 2.0 mM Trolox, 0.026 units/$\mu$L pyruvate kinase/lactic dehydrogenase, and 26.6 mM phosphoenolpyruvic acid.
We note that the sample is sensitive to the buffer pH and mixture incubation time. For our experimental conditions, the mixture pH is around 6.4 and we perform the experiments within 2 hours of constructing the mixture.
II.5 Microscope setup
We conducted the experiments using an automated widefield epifluorescence microscope (Nikon Ti-2), custom-modified for two additional imaging modes: epi-illuminated pattern projection and LED-gated transmitted light. Light patterns from a programmable DLP chip (EKB Technologies DLP LightCrafter™ E4500 MKII™ Fiber Couple) were projected onto the sample via a user-modified epi-illumination attachment (Nikon T-FL). The DLP chip was illuminated by a fiber-coupled 470 nm LED (ThorLabs M470L3). The epi-illumination attachment featured two light-path entry ports: one for the projected pattern light path and the other for a standard widefield epi-fluorescence light path. These light paths were combined using a dichroic mirror (Semrock BLP01-488R-25). The magnification of the epi-illumination system was calibrated to ensure the camera’s imaging sensor (FliR BFLY-U3-23S6M-C) was fully illuminated when the entire DLP chip was activated. Micro-Manager software, running custom scripts, controlled the pattern projection and stage movement. For the transmitted light path, we replaced the standard white-light brightfield source (Nikon T-DH) with an electronically time-gated 660 nm LED (ThorLabs M660L4-C5) to minimize light-induced dimerization during brightfield imaging.
II.6 Light pattern construction
In this work, we followed the illumination protocol from Ref. [6], using pulsed light to activate motor-linking. A light pattern is projected onto the active matter system every 10 seconds, with each pulse lasting 30 milliseconds. This protocol ensures sufficient motor activation for network contraction and sustained fluid flow. Shorter illumination periods fail to generate enough cross-linked microtubules, while longer or continuous illumination causes motor activation outside the intended region, disrupting the network and fluid flow.
For generating V-shape light patterns, we utilized a Python script to calculate the V shape’s coordinates and mapped them onto a 1280 by 800 pixel canvas. The pattern was displayed as white on a black background. The final V shape pattern was saved as a TIFF file at the specified location for subsequent experimental use. A custom Java script was used to read the tiff file and project the custom light pattern to the active matter system in micro-manager.
II.7 Image processing for centerline-curvature measurement
We first separated the left and right arms of the networks. We used the “skeletonize” function from the scikit-image package [18] in Python to extract the skeleton of each arm. We then used the FilFinder package [19] to pick out the longest branch of the skeleton, which is the centerline of the arm.
To measure the curvature of the centerline, we first fitted the centerline with a parabola $y=ax^{2}+bx+c$. The local curvature $\kappa$ of the centerline is is $\kappa=y^{\prime\prime}/(1+y^{\prime 2})^{3/2}$ and the average curvature $[\kappa]$ is
$[\kappa]=\frac{\int_{x_{0}}^{x_{t}}y^{\prime\prime}/(1+y^{\prime 2})\mathrm{d}x}{\int_{x_{0}}^{x_{t}}(1+y^{\prime 2})^{1/2}\mathrm{d}x},$ (S36)
which can be numerically integrated using $y^{\prime\prime}=2a$ and $y^{\prime}=2ax+b$. $x_{0}$ and $x_{t}$ are the x-coordinates of the starting and ending points of the centerline, respectively.